Viu 1.u For u = v _ -3 1, find viu, l|v||, 4, and a unit vector in the 2 1-1) direction opposite of u.
Answer #1

Gerwen u l .-1-5 = (1, -3, 2,-1) & u(-2,1,3,2) Voll = 1x (-2) + (-3) X 2 + 2x3 + (-1) x2 - -2-3 + 6-2 - - 7+6 = -1 TV. u=-1 IHoll= (-2,1,3,2). (2,1,3,2) - (-2) × (-2) + 1x2 + 3x3 + 2x2 - 4 + 1 + 9 + 4 = 16 vou Unit vector of u = ů = u 11W ilul = √ (-